DPS Calculation Thress Industries just paid a dividend of $3.00 a share (i.e., D0 = $3.00). The dividend is expected to grow 9% a year for the next 3 years and then 15% a year thereafter. What is the expected dividend per share for each of the next 5 years? Round your answers to the nearest cent.

  1. D1 = $  

  2. D2 = $  

  3. D3 = $  

  4. D4 = $  

  5. D5 = $  
Answer #1

(girem) & gi= 0.0 9 & 92= 0·15 $3.00 (1+ 0·09) = $3.00 (1+ o.09)2=$3.56 $ 3.00(1+0.09)3 Given:- Do $3.00 Do x(1 +9,) = Do x